Sketch the gradient field ∇f of the following functions:
f(x , y) = (x^2 −y^2)/4
1
Expert's answer
2020-10-15T17:16:22-0400

"\bigtriangledown (f) =<f_{x},f_{y}>=<\dfrac{x}{2},-\dfrac{y}{2}>" Hence for sketching, the vector with initial point "(x,y)"points parallel to the vector "(\frac{x}{2},-\frac{y}{2})." So the gradient looks like